Suppose that the national average for the math portion of the College Boards SAT is 548. The College Board periodically rescales the test scores such that the standard deviation is approximately 100. Answer the following questions using a bell-shaped distribution and the empirical rule for the math test scores. If required, round your answers to two decimal places. (a) What percentage of students have an SAT math score greater than 648? (b) What percentage of students have an SAT math score greater than 748? (c) What percentage of students have an SAT math score between 448 and 548? (d) What is the z-score for student with an SAT math score of 635? (e) What is the z-score for a student with an SAT math score of 425?
Answer #1

Here we have : mu = 548, sigma = 100

a) Here we need to find

p ( x > 648 )

Z一,1 . 648-548

= p ( z > 1 )

= 1 - p ( z leq 1 )

= 1 - 0.8413

= 0.1587

So percentage is 15.87 %.

b)  

Here we need to find

p ( x > 748 )

2.-μ 748-548

= p ( z > 2 )

= 1 - p ( z leq 2 )

= 1 - 0.9773

= 0.0227

So percentage is 2.27 %.

c)

Here we need to find

p ( 448 < x < 548 )

448-548 x_μ 548-548 100

= p ( -1 < z < 0 )

= p ( z < 0 ) - p ( z < -1 )

= 0.5 - 0.1587

= 0.3413

So percentage is 34.13 %.

d) Z score for 635 .

635-548 100 x-μ 0.87

e)

Z score for 425 .

x-μ 425-548 100 1.23